OpenStudy (anonymous):

A population P obeys the logistic model. It satisfies the equation dP/dt=(3/700)P(7−P) for P>0 (a)The population is increasing when P= (b) The population is decreasing when P= (c) Assume that P(0)=3 Find P(60) P(60)=

OpenStudy (anonymous):

breaking it up i would assume that it comes out to this if im not mistaken\[\int\limits \frac{ 700 }{ 3[P(7-P)] }dP=\int\limits 1dt\]

OpenStudy (anonymous):

Yes you are right sjerman

OpenStudy (anonymous):

what would i do for integrating that, should i expand the denominator first then try for u-sub or integration by parts? maybe just quotient rule?

OpenStudy (anonymous):

you have to do partial fraction decomposition and then integrate it , that makes it very simple

OpenStudy (anonymous):

do u know how to do partial fraction decomposition?

OpenStudy (anonymous):